what is the maximum value of 8-2x-x^2
Answer by Alan3354(69443)

what is the maximum value of 8-2x-x^2
----------
f(x) = 8-2x-x^2
That's a parabola -x^2 - 2x + 8 in ax^2 + bx + c form
The max is the vertex at x = -b/2a
x = 2/-2 = -1
f(-1) = -1 +2 +8 = 9